OpenStudy (anonymous):

Steve is adding wallpaper to a living room wall and he needs to know how much wallpaper to buy. If the wall is 8.5 feet tall and 12.5 wide, how much wallpaper should he buy?

OpenStudy (anonymous):

help!

OpenStudy (mendicant_bias):

Hey. Is there a price given for square footage of wallpaer, or is it just asking for the amount of wallpaper by itself? (Don't mean to ask you again, but just making entirely sure.)

OpenStudy (anonymous):

the wallpaper by its self.

OpenStudy (mendicant_bias):



OpenStudy (mendicant_bias):

Sorry, computer issues. One moment.

OpenStudy (anonymous):

that is okay

OpenStudy (mendicant_bias):

As a general rule, for any rectangle, the area will always be the product of any given side by the side perpendicular to it. The reason i'm not using labels like "length", "width", or "height" to describe any of those sides is that they're totally up to you to call them that. Any object can have any dimension along it called length, width, or height, depending on what people feel like labeling it. So, in this situation: \[A = w*h, --- --- --- w = 12.5, a = 8.5\] \[A = w*h = (12.5)(8.5)\] Does this make sense?

OpenStudy (mendicant_bias):

The number that would come out of it would be the area, in units of "square feet". So, you could say that Steve needs to purchase (w*h) square feet of wallpaper. Square feet can also be abbreviated as "sq. ft."

OpenStudy (anonymous):

so i would have to solve 8.5=12.5*h

OpenStudy (mendicant_bias):

Wait, what? Sorry. The given height is 8.5, and the given width is 12.5, correct. If \[A = w*h \] And here, \[w = 12.5\] and \[ h = 8.5\] So what would that equation look like if you substituted in w and h?

OpenStudy (anonymous):

then my answer would be 106.25 right?

OpenStudy (mendicant_bias):

Yeah, but let's make sure this is all right. In your last post, you made the equation look like you were solving for h, when h is known. In the equation given, A represents area, which is what you're looking for, right? So what do you think it would look like? If I have\[A = w*h\]What would that look like where A is being solved for and \[w = 12.5, h = 8.5\]?

OpenStudy (anonymous):

a=12.5*8.5

OpenStudy (mendicant_bias):

Yes! Exactly! And yeah, your answer would be 106.25. But wait a minute, do you know what "significant figures" are?